Porter governor has equal arms each 315 mm long and pivoted on the axis of
rotation. Each ball has a mass of 7.3 kg and the mass of the central load on the
sleeve is 28 kg. The radius of rotation of the ball is 175 mm when the governor
begins to lift and 220 mm when the governor is at maximum speed. Compute the
minimum and maximum speed and range of speed of the governor when the arm
length is increased by 33%. Discuss the effect of speed range when maximum
radius of rotation is equal to the arm length and also justify the usage of variable
arm length in the governor.
Answer #1

Criven Poxter Governor arm length = 315 mm=0.315m moss of ball (m) = 7.3kg moss of central load (M) = 88kg 8,= 175 mm= 0.175mMinimum speed when 8 = 0.175 m In A PIGB in fig 0, the - .Jpg?- Ba? -J0.4193-(0.175)2 hi = 0.38m m + 395 M X 2 Ni 5 73 +28 89N = mtx on see m Na = \o 2 x m. range of speed a No-N.. = 109-106 = 3 8.p.m

So after the increase in arm length by 33%,

Minimum speed = 106 r.p.m.

Maximum speed = 109 r.p.m.

Range of speed = N2 - N1 = 109 - 106 = 3 r.p.m.

If the radius of rotation is equal to the arm length, it means the angle between then is zero. This is never can happen.
Both are equal means they overlap with each other.

If you have any quarry related to this then comment below.